# EFRame: Deeper Reasoning via Exploration-Filter-Replay Reinforcement Learning Framework

### v3

@@ Abstract
Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs). Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks. To address these limitations, we introduce EFRame, an Exploration-Filter-Replay framework that systematically augments GRPO along three critical dimensions. EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples. EFRame establishes a complete and stable learning cycle, guiding the model through a structured transition from exploration to convergence. Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO. Furthermore, EFRame not only enables fine-grained categorization of training samples for deeper insight into their contributions, but also introduces an efficient and precise mechanism for entropy control, which is critical for balancing exploration and convergence in RL training. Our code is available at https://github.com/597358816/EFRame.
@@ 1 Introduction
RLVR relies on extracting reward signals from the model’s own pass@$k$ outputs to improve its pass@1 performance. However, this design inherently couples final performance to the choice of $k$, as the reward signal vanishes when all sampled responses are either correct or incorrect, making it impossible for the model to extract useful knowledge from such data. This constrains RLVR’s learning capacity to the base model’s pass@$k$, limiting further exploration and improvement Yue et al. (2025). Although increasing the sampling number $k$ per prompt may seem like a straightforward solution, it typically yields only marginal improvements while significantly reducing sample efficiency Lin et al. (2025). This results in GRPO exhibiting limited exploration, and low sample efficiency, and indirectly leads to training instability in the later stages due to the disappearance of reward signals Yu et al. (2025); Li et al. (2025).
To address the aforementioned issues, we propose EFRame: an Exploration-Filter-Replay RL Framework. For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout phase to enable targeted exploration. During this process, the model generates a small number of high-quality positive samples along with a large number of low-quality negative samples. To reduce training variance and improve efficiency, we apply an online filtering mechanism to discard low-quality samples. Meanwhile, high-advantage samples are stored in a replay buffer and reused during training to amplify their impact and enhance model stability.
@@
    Precise and Efficient Entropy Control: EFRame offers a practical and effective approach to entropy control by jointly tuning the additional rollout temperature and the number of replayed samples. This mechanism provides more flexible and stable entropy regulation than traditional reward-based entropy regularization, supporting both efficient exploration and stable convergence.
EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning. By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models. Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm. Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
@@ 2 Related Work
The early application of RL in LLMs and MLLMs OpenAI (2023); Team et al. (2024); Zhu et al. (2023); Wei et al. (2023); Liu et al. (2023) was primarily reflected in RLHF, where algorithms such as PPO were employed to align language model outputs with human preferences and Direct Preference Optimization (DPO) Rafailov et al. (2024) bypass the need for reward modeling and instead optimize the model policy directly from preference data. Recently, RL has advanced toward integrating rule-based reward mechanisms with sophisticated optimization techniques, as exemplified by models such as DeepSeek-R1 Guo et al. (2025) and Kimi-1.5 Team et al. (2025). As a lightweight variant of PPO, GRPO Shao et al. (2024) significantly improves training efficiency, but at the cost of increased instability. A series of subsequent studies have introduced optimizations from GRPO. These include gradient loss refinements, as demonstrated in DAPO Yu et al. (2025) and Dr.GRPO Liu et al. (2025); online filtering methods like Meng et al. (2025); advantage enhancement techniques, exemplified by VL-Rethinker Wang et al. (2025a) and Luffy Yan et al. (2025); clip acceleration methods such as CPPO Lin et al. (2025); curriculum learning strategies explored in Chen et al. (2025b; c); and unsupervised RL such as Zhang et al. (2025); Wei et al. (2025a); Zhao et al. (2025); Agarwal et al. (2025). While existing methods typically focus on individual components, we systematically review the fundamental limitations of GRPO and propose a comprehensive framework with an integrated solution that addresses three critical aspects: exploration capacity, training stability, and sample efficiency. This helps create a more robust and effective GRPO-based training paradigm.
3 Overview of GRPO
3.1 Gradient Loss
GRPO has been employed to enhance the reasoning capabilities of large language models, particularly in mathematical problem-solving. Compared to PPO, GRPO improves computational efficiency by eliminating the need for a separate critic model. Instead, it estimates the advantage directly using group-normalized rewards. Given a question $q$ and a corresponding verifiable answer $a$, GRPO generates a set of responses $O=\{o_{i}\}_{i=1}^{G}$ using the policy model $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{(q,a)\sim\mathcal{D},\ O=\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}(\cdot|q)}}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t},\ \operatorname{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]-\beta D_{KL}[\pi_{\theta}||\pi_{ref}]\right\},\end{split}$ (1)
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i}<t)}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i}<t)}$ and $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}$.
3.2 Limitations of GRPO
Although GRPO has demonstrated strong performance in various settings, it encounters several practical challenges when applied to complex reasoning tasks. In the following discussion, we highlight three main issues: limited exploration capacity, training collapse, and low sample efficiency. These limitations reveal the need for more robust and adaptive RL strategies.
Limited Exploration Capacity. GRPO lacks effective mechanisms to promote exploration, leading to persistently low entropy levels in the early stages of training and an increased risk of convergence to local optima. Moreover, it does not differentiate between prompts of varying difficulty, which limits its ability to stimulate deeper reasoning capabilities. As a consequence, only superficial knowledge is activated, while deeper knowledge remains underutilized.
((a))
((b))
((c))
Figure 1: Increasing the sampling number to $G=20$ provides a performance improvement over $G=5$, but it incurs a fourfold increase in per-step runtime, leading to very low sample efficiency in practice.
Low Sample Efficiency. GRPO computes sample advantage based on relative advantage. When the number of rollouts is small, it suffers from a severe issue of advantage vanishing, leading to a lack of effective training signals. However, as the number of samples increases, the relative advantage becomes diluted, making it harder for high-quality samples to exert sufficient influence. This results in low sample efficiency: although the computational overhead grows nearly linearly with the number of samples, the corresponding performance improvement is significantly diminished, as shown in Fig. 1.
Instability. During the later stages of training, the GRPO algorithm may exhibit signs of instability or collapse, characterized by a sharp decline in reward and sudden spikes in both gradient and policy entropy, as illustrated in Fig. 2. In extreme cases, this can escalate into a gradient explosion. Such instability presents a serious challenge to reliable long-horizon optimization, particularly in complex reasoning tasks where consistent gradient signals are critical for convergence. One contributing factor is the increased variance caused by the disappearance of relative advantages in the later stages of training. Supplementing with additional samples with non-zero advantages can help alleviate this issue.
((a))
((b))
((c))
Figure 2: When training Qwen2.5-VL-7B-Instruct on the Geo3k dataset using the GRPO algorithm, a collapse occurred.
@@ 4 Methodology
To address the issues in the previous section, we propose the EFRame, which consists of three main components: additional rollout, online filtering and replay buffer, as illustrated in Figure 3.
Figure 3: The overall workflow of EFRame builds upon the vanilla GRPO framework by introducing three key components: additional rollout, online filtering, and replay buffer.
4.1 Regular Rollout
For each batch of prompts, we begin with a regular rollout phase, following the original GRPO training process. This initial rollout serves two key purposes:
@@ 4.1 Regular Rollout
  • Identify more challenging prompts that fail to yield positive samples, which will be handled in subsequent stages.
Valuable samples are temporarily stored in the buffer for gradient updates.
@@ 4.2 Additional Rollout
In GRPO-based algorithms with binary rewards (Premise, omitted in the following), difficult prompts are defined as those for which all sampled responses within the initial $G_{1}$ rollouts ($G_{1}$ is a smaller integer, such as 5) fail to obtain a positive reward, where $\bar{\mathcal{D}}\subseteq\mathcal{D}$ and:
$\bar{\mathcal{D}}=\left\{(q,a)~{}:~{}|\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}~{}:~{}\text{is\_equivalent}(a,o_{i})\right\}_{i=1}^{G_{1}}|=0\right\},$
@@ 4.3 Online filter
While additional rollout enhances the model’s ability to explore complex prompts, it also introduces computational overhead and potential training noise. In addition to filtering overlong samples, we design two additional filtering strategies to further enhance the stability and efficiency of the training process.
@@
$O_{H}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}>0\right\};$
@@
$O_{L}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}<0\right\}.$
@@
[…]
$\qquad\qquad\quad{}-\,\beta\,D_{\mathrm{KL}}\bigl{[}\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\bigr{]}\Bigr{)}\Biggr{]}.$
To prevent token-length bias, we compute the policy loss using the token-mean aggregation strategy.
4.4 Replay buffer
@@
The high-quality samples identified through additional rollouts are typically responses to difficult prompts and thus are with relatively low probability. Using them only once yields limited and short-lived gradient effects, leading to an increase in model entropy. By incorporating these samples into the replay buffer, their probability of reoccurrence is effectively increased during training, enabling the model to transition from entropy increase to eventual convergence. In fact, the replay buffer here serves to reduce policy entropy; detailed discussions can be found in the following discussion section.
4.5 Discussion
The above subsections yield several practical insights regarding the interplay between Exploration-Filter-Replay:
Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high-temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$. According to Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing the collapse of the entropy and maintaining effective exploration. At the same time, an excessive presence of low-quality samples can cause a nosedive in entropy, which is difficult to control. To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
Replay Buffer: By storing high-quality positive samples within the replay buffer, the model can repeatedly utilize their gradient information for subsequent updates. As the model has already been updated on these (problem, answer) pairs $(x,y)$, the recalculated $\log\pi_{\theta}(y\mid x)$ becomes larger upon rehearing these pairs. According to the theorem, this shift leads to a reduction in policy entropy, preventing excessive growth in entropy.
@@ 5 Experiments
To validate the effectiveness of our methods, we demonstrate our experiment setup and the results in the following sections.
Table 1: Comparative results across multiple multimodal reasoning benchmarks indicate that our method achieves excellent performance.
Models Size MathVision MathVerse MathVista We-Math Average
Close-Source Models
GPT-4o Hurst et al. (2024) - 33.95 48.83 59.50 65.00 51.82
Claude-3.5-Sonnet Anthropic (2024) - 46.48 57.64 68.20 73.05 61.34
Open-Source Multi-Modal Large Reasoning Models
QvQ-Preview Bai et al. (2025) 72B 35.56 52.81 69.20 65.29 55.72
Skywork R1V Peng et al. (2025) 38B 39.31 40.91 67.50 60.06 51.95
Open-Source 7B Multi-Modal Models
MM-Eureka Meng et al. (2025) 7B 28.36 49.52 70.40 66.03 53.58
OpenVLThinker Deng et al. (2025) 7B 25.30 42.79 64.10 63.91 49.03
ThinkLite-VL Wang et al. (2025b) 7B 27.71 47.41 69.00 67.99 53.03
VLAA-Thinker Chen et al. (2025a) 7B 26.61 47.49 66.20 66.49 51.70
Qwen2.5-VL-7B-Instruct with RL
Qwen2.5-VL Bai et al. (2025) 7B 24.87 43.83 66.30 62.87 49.47
Qwen2.5-VL+GRPO 7B 29.11 47.51 72.60 67.53 54.19
Qwen2.5-VL+DAPO 7B 27.92 48.48 72.30 69.08 54.45
Ours 7B 28.82 (+3.95) 48.81 (+4.98) 73.40 (+7.10) 69.48 (+6.61) 55.13 (+5.66)

### v5

# inkscapearea=page

@@ Abstract
Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs). Group Relative Policy Optimization (GRPO), a lightweight variant of Proximal Policy Optimization (PPO), improves efficiency but suffers from limited exploration and training instability, limiting its effectiveness on complex reasoning tasks. To address these challenges, we introduce EFRame, an Exploration-Filter-Replay framework that augments GRPO across three dimensions: additional rollouts enable deeper and more targeted exploration, online filtering removes low-quality samples to stabilize gradients and accelerate training, and experience replay amplifies rare yet informative trajectories for stable convergence. This unified framework establishes a principled training cycle that balances exploration, efficiency, and stability. Experiments on diverse reasoning benchmarks demonstrate that EFRame achieves consistent gains, including a 37.9% relative improvement on Geometry3K over GRPO. EFRame further supports fine-grained sample categorization and precise entropy control, highlighting it as a robust solution for advancing deeper reasoning in LLMs.
@@ 1 Introduction
Despite its efficiency, GRPO suffers from limited exploration and training instability on complex reasoning tasks. We conducted extensive experiments and categorized the failed GRPO training cases into two types, as illustrated in Fig. 1, which shows representative runs on Qwen2.5-VL-7B-Instruct with the Geometry3K dataset. In the first example (GRPO-1 in Fig. 1), exploration was severely limited: the policy entropy continuously decreased from an initial value of 0.25 throughout training, making it difficult for the model to perform effective exploration and resulting in insufficient fine-tuning improvements. In the second example (GRPO-2 in Fig. 1), the training suffered an irreversible collapse, marked by an arch-shaped fall of rewards and accuracy during mid-training, followed by a breakdown of model performance that could not be recovered.
((a))
((b))
((c))
Figure 1: Two prominent issues of GRPO: limited exploration and training instability, when training on Qwen2.5-VL-7B-Instruct with the Geometry3K dataset.
GRPO derives reward signals from relative advantages computed within rollout groups. This design inherently ties the training process to the accuracy of groups Yue et al. (2025). When all sampled responses are either correct or incorrect, the relative advantage collapses to zero, yielding no effective learning signal. In long-horizon training, this advantage-vanishing effect amplifies gradient variance, which often manifests as gradient explosions and continuous reward degradation, eventually leading to irreversible training collapse Yu et al. (2025); Li et al. (2025). Collectively, these issues constrain GRPO’s ability to sustain exploration, preserve model accuracy, and maintain stability on challenging reasoning tasks.
To address the aforementioned issues, we propose EFRame: an Exploration-Filter-Replay RL Framework. For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout to enable targeted exploration. During this process, the model generates a small number of high-quality positive samples along with a large number of low-quality negative samples. To reduce training variance and improve efficiency, we apply an online filtering mechanism to discard low-quality samples. Meanwhile, high-advantage samples are stored for experience replay to amplify their impact and enhance model stability.
@@
    Efficient Entropy Control: EFRame provides a practical and effective method for entropy control by jointly tuning the additional rollout temperature and the number of replayed samples. Compared to traditional reward-based entropy regularization, this mechanism enables more flexible and stable entropy regulation, thereby facilitating both efficient exploration and stable convergence.
Overall, EFRame provides a systematic solution to the inherent trade-offs among stability, efficiency, and exploration capacity in RL. By effectively identifying and leveraging deeper latent knowledge, it enhances the expressiveness and learning potential of LLMs. Through the coordinated integration of targeted exploration, online filtering, and experience replay, EFRame not only addresses the limitations of GRPO but also establishes a principled and controllable training paradigm. Empirical results confirm its efficacy: on the challenging Geometry3K benchmark, EFRame outperforms GRPO by 37.9%, and it achieves consistent improvements across diverse mathematical and multimodal reasoning benchmarks, demonstrating strong robustness and generalizability.
EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in RL. By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models. Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm. Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 37.9%, and it achieves consistent gains across a wide range of mathematical and multimodal reasoning benchmarks, demonstrating strong generalizability and robustness.
@@ 2 Related Work
The early application of RL in LLMs and MLLMs OpenAI (2023); Team et al. (2024); Zhu et al. (2023); Wei et al. (2023); Liu et al. (2023) was primarily reflected in RLHF, where algorithms such as PPO were employed to align language model outputs with human preferences and Direct Preference Optimization (DPO) Rafailov et al. (2024) bypass the need for reward modeling and instead optimize the model policy directly from preference data. Recently, RL has advanced toward integrating rule-based reward mechanisms with sophisticated optimization techniques, as exemplified by models such as DeepSeek-R1 Guo et al. (2025) and Kimi-1.5 Team et al. (2025). As a lightweight variant of PPO, GRPO Shao et al. (2024) significantly improves training efficiency, but at the cost of limited exploration and training instability.
A series of subsequent studies have introduced optimizations from GRPO. DAPO Yu et al. (2025) and Dr.GRPO Liu et al. (2025) adjust the policy gradient from a more global perspective, but they do not directly address the aforementioned issues. Entropy-driven exploration methods, e.g., Cheng et al. (2025); Chen et al. (2025), introduce entropy regularization into reward or advantage functions to encourage exploration. However, this strategy only provides a coarse trend, making entropy difficult to control precisely and stably, while also introducing biases that are hard to estimate. Instead, we adopt an exploration method with limited bias—high quality sampling—which enables more effective exploration with controlled deviation. Online filtering accelerations, such as Meng et al. (2025); Lin et al. (2025), further strengthen the efficiency advantage of GRPO, but their contribution to training effectiveness is almost negligible. Experience replay approaches, including VL-Rethinker Wang et al. (2025), RLEP Zhang et al. (2025), and RePO Yan et al. (2025), enhance GRPO from the perspective of convergence stability. However, experience replay may also aggravate the entropy collapse problem in GRPO, and indiscriminate use can hinder the model’s ability to explore more deeply. EFRame addresses these challenges by encouraging exploration through conditional sampling, accelerating training with online filtering, and achieving stable convergence via experience replay, thereby delivering consistent improvements to GRPO in exploration, stability, and convergence.
Figure 2: The overall workflow of EFRame builds upon GRPO by introducing three key components: additional rollout, online filter, and experience replay.
@@ 3 Methodology
To resolve the limited exploration and training instability in GRPO, we propose the EFRame, which consists of three main components: additional rollout, online filtering, and experience replay, as illustrated in Figure 2.
3.1 GRPO and Regular Rollout
Group Relative Policy Optimization (GRPO) directly estimates the advantage using group-normalized rewards. Given a question $q$ and a corresponding verifiable answer $a$, GRPO generates a set of responses $O=\{o_{i}\}_{i=1}^{G}$ using the policy model $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{(q,a)\sim\mathcal{D},\ O=\{o_{i}\}_{i=1}^{G_{1}}\sim\pi_{\theta_{old}(\cdot|q)}}\\ &\frac{1}{G_{1}}\sum_{i=1}^{G_{1}}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t},\ \text{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]-\beta D_{KL}[\pi_{\theta}||\pi_{ref}]\right\},\end{split}$ (1)
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i}<t)}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i}<t)}$ and $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{1}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{1}})}$.
In EFRame, training begins with a regular rollout phase following this GRPO formulation. This initial rollout serves two key purposes:
@@ 3.1 GRPO and Regular Rollout
  • Identify more challenging prompts that fail to yield positive samples, which will be addressed in subsequent stages.
@@ 3.2 Additional Rollout
In GRPO-based algorithms with binary rewards (Premise, omitted in the following), difficult prompts are defined as those for which all sampled responses within the initial $G_{1}$ rollouts ($G_{1}$ is a small integer, such as 5) fail to obtain a positive reward, where $\bar{\mathcal{D}}\subseteq\mathcal{D}$ and:
$\bar{\mathcal{D}}=\left\{(q,a)~:~|\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}~:~\text{is\_equivalent}(a,o_{i})\right\}_{i=1}^{G_{1}}|=0\right\}.$
@@ 3.3 Online filter
While additional rollout enhances the model’s ability to explore complex prompts, it also introduces computational overhead and potential training noise. In addition to filtering overlong samples introduced in DAPO Yu et al. (2025), we design two additional filtering strategies to further enhance the stability and efficiency of the training process.
@@
$O_{H}=\left\{\{o_{i}\}\sim\pi^{t_{a}}_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}>0\right\};$
@@
$O_{L}=\left\{\{o_{i}\}\sim\pi^{t_{a}}_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}<0\right\},$
where $\pi^{t_{a}}_{\theta_{old}}$ denotes the distribution of $\pi_{\theta_{old}}$ under temperature $t_{a}$. In our setting, $t_{a}=1.2$ by default.
@@
$\mathcal{J}(\theta)$ $=\mathbb{E}\Biggl[\frac{1}{\displaystyle\sum_{o_{i}\in O_{R}\cup O_{H}}\lvert o_{i}\rvert}\sum_{o_{i}\in O_{R}\cup O_{H}}\sum_{t=1}^{\lvert o_{i}\rvert}\Bigl(\min\bigl\{\,r_{i,t}(\theta)\,\hat{A}_{i,t},\;\mathrm{clip}\bigl(r_{i,t}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}}\bigr)\,\hat{A}_{i,t}\bigr\}\Biggr].$ (2)
To prevent token-length bias, we compute the policy loss using a token-mean aggregation strategy. Furthermore, following the conclusion of DAPO, and given that catastrophic forgetting is relatively minor in RL Lai et al. (2025), we omit KL regularization during training to allow a more thorough optimization of the model.
3.4 Experience replay
@@
The high-quality samples identified through additional rollouts are typically responses to difficult prompts and thus are with relatively low probability. Using them only once yields limited and short-lived gradient effects, leading to an increase in model entropy. By incorporating these samples into the replay buffer, their probability of reoccurrence is effectively increased during training, enabling the model to transition from entropy increase to eventual convergence. In fact, the experience replay here serves to reduce policy entropy; detailed discussions can be found in the following subsection.
@@ 4 Experiments
To validate the effectiveness of our methods, we present the experimental setup and results in the following sections. Our work is based on the EasyR1 and VeRL frameworks Yaowei et al. (2025); Sheng et al. (2025), and we compare with the RL baselines GRPO Shao et al. (2024) and its improved variant DAPO Yu et al. (2025). See implementation details in Appendix A.
Table 1: Comparative results across text-only (LLM) and multimodal (VLM) reasoning benchmarks.
Text Benchmarks AIME’24 MATH GSM8K Minerva Average
Qwen2.5-math-7B
Yang et al. (2024)
13.3 65.5 65.4 11.0 38.8
GRPO Shao et al. (2024) 26.7 77.6 87.1 29.0 55.1
DAPO Yu et al. (2025) 30.0 78.3 87.6 34.2 57.5
EFRame (ours) 36.7 (+23.4) 80.4 (+19.9) 87.9 (+22.5) 35.7 (+24.7) 60.2 (+21.4)
Multimodal Benchmarks MathVision MathVerse MathVista We-Math Average
Qwen2.5-VL-7B-Instruct
Bai et al. (2025)
24.87 43.83 66.30 62.87 49.47
GRPO Shao et al. (2024) 29.11 47.51 72.60 67.53 54.19
DAPO Yu et al. (2025) 27.92 48.48 72.30 69.08 54.45
EFRame (ours) 28.82 (+3.95) 48.81 (+4.98) 73.40 (+7.10) 69.48 (+6.61) 55.13 (+5.66)